OpenStudy (anonymous):

Tough one! For what values of a and b is the following equation true? limx->0 ((sin2x)/x^3)+a+(b/x^2)) = 0 @Calculus1

OpenStudy (amistre64):

it would have to amount to limit of the sums maybe lim (sin2x)/x^3 lim a lim b/x^2

OpenStudy (anonymous):

Exactly! Good, I was on the right track. Okay so: I figured lim sin2x/x^3 = 2/0 lim a=a lim b/x^2= 0/2

OpenStudy (amistre64):

the limit of a = a .... so when does the lim of b/x^2 + lim sin(2x)/x^3 = -a

OpenStudy (anonymous):

Wait, is that a rhetorical question? Was I wrong? Is the limit of a not a?

OpenStudy (amistre64):

the easiest thing to do would be to get a limit for sin(2x)/x^3 to calibrate this stuff with

OpenStudy (amistre64):

the limit of a constant is the constant

OpenStudy (amistre64):

the whole stuff has to equal 0, so the other stuff has to amount to -a

OpenStudy (amistre64):

a - a = 0

OpenStudy (anonymous):

So Do i just plug in the limits I found for the other two parts of the equation?

OpenStudy (anonymous):

aka 2/0 and 0/2, respectively

OpenStudy (amistre64):

h = sin(2x)/x^3 h' = 2cos(2x)/3x^2 h'' = -4sin(2x)/6x h''' = -8cos(2x)/6 lim h''' = -8 cos(0)/6 = -4/3 right?

OpenStudy (amistre64):

ugh .... thought lhopital would do it; but the wolf says that goes up to infinity

OpenStudy (anonymous):

lol, "the wolf"

OpenStudy (amistre64):

you sure you got it typed up right? http://www.wolframalpha.com/input/?i=solve+for+a+and+b+lim_%28x+to+0%29+%28%28sin2x%29%2Fx%5E3%29%2Ba%2B%28b%2Fx%5E2%29%29+%3D+0 it says no solutions

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Look what I found though.... Lol yahoo answers: http://answers.yahoo.com/question/index?qid=20091110070715AAcJ0Vq

OpenStudy (anonymous):

Looks like you were right with -4/3. How come the wolf is saying that is wrong though..

OpenStudy (amistre64):

thats where my lhopital went bad, i kept at it till the bottom was clear ....

OpenStudy (anonymous):

Oh, that isn't allowed?

OpenStudy (amistre64):

apparently not, its not in the correct indeterminate form to reapply it

OpenStudy (anonymous):

Any ideas or are you stumped?

OpenStudy (amistre64):

the wolf says those solutions from yahoo are false

OpenStudy (amistre64):

the answer is sometimes: does not exist ....

OpenStudy (anonymous):

Lets see if agreene can enlighten us?

OpenStudy (amistre64):

since the limit fails to exist on the sin part, there is nothing you can do

OpenStudy (agreene):

Since: \[\lim_{x \rightarrow 0} \frac{\sin (2x)}{x^3}= \infty\] I should think that in order for this to approach 0... a and/or b must be approaching negative infinity at the same rate(s).

OpenStudy (amistre64):

a cant change since its constant

OpenStudy (amistre64):

unless its 0

OpenStudy (amistre64):

but then infinity - infinity is not defined to be zero

OpenStudy (agreene):

indeed, thats why I said it has to approach -infty at the same rate (to perpetually cancel the movement towards infinity) I don't know of any numbers you can plug in to make this limit 0...

OpenStudy (amistre64):

well, a=0 and b=-1 seems to work

OpenStudy (agreene):

Oh i'm dumb sorry, I used pos 1.

OpenStudy (agreene):

O_o didnt change anything with -1 http://www.wolframalpha.com/input/?i=limx-%3E0+%28%28sin2x%29%2Fx%5E3%29%2B0-%281%2Fx%5E2%29%29+

OpenStudy (amistre64):

i spose inf-inf doesnt equal 0, unless when it does ....

OpenStudy (amistre64):

wrap your equation

OpenStudy (amistre64):

at the moment its taking lim sin and adding it to the rest

OpenStudy (agreene):

wrapping doesnt make a difference... I dont know why it's calling it 0 for you and infty for me O_o http://www.wolframalpha.com/input/?i=limx-%3E0+%28%28%28sin2x%29%2Fx%5E3%29%2B0-%281%2Fx%5E2%29%29

OpenStudy (amistre64):

lol .... i asked it it = 0 and it says true; then shows steps. the wolf has gone schitzoid i spose

OpenStudy (anonymous):

NOOOOOOOOOOOOOOOOOOOO

OpenStudy (agreene):

I see why, you used sin(x) It's sin(2x)... sin(x) makes it go to 0 sin(2x) makes it go to infty.

OpenStudy (amistre64):

i just noticed that too ....

OpenStudy (amistre64):

fine. a=0, b=-2 lol

OpenStudy (amistre64):

when i leave off the =0 , it says the limit is -4/3 lol

OpenStudy (agreene):

To answer the question, I think: \[\forall a \lim f(x) \neq 0\] and \[\forall a \lim f(x) \neq 0\] Even though I dont have any proof, lol

OpenStudy (agreene):

2nd one is supposed to be forall b

OpenStudy (anonymous):

what does that V symbol mean? Is that an angle symbol?

OpenStudy (agreene):

Its an upside down A... it means For all (variable) there should be a :: after, but I forgot those. Math shorthand is useful... instead of writing: For all a such that limit of the function doesnt equal 0 you can just write: \[\forall a :: \lim f(x) \neq 0\]

OpenStudy (anonymous):

Wait so that means, "for all a such that the limit of the function doesn't equal 0", okay

OpenStudy (agreene):

Some people prefer | instead of :: for such that, I was taught with :: but most people use | nowdays.. meh

OpenStudy (agreene):

They come in handy: http://en.wikipedia.org/wiki/Table_of_mathematical_symbols

OpenStudy (agreene):

It basically says, there is no A or B that exist by that will make it equal 0.

OpenStudy (amistre64):

if a=0 and b=-2 we can lhopital this into submission i believe to get \[\frac{sin(2x)}{x^3}-\frac{2}{x^2}\] \[\frac{x^2sin(2x)-2x^3}{x^5},\frac{0}{0}\] \[\frac{2xsin(2x)+2x^2 cos(2x)-6x^2}{5x^4},\frac{0}{0}\] \[\frac{2sin(2x)+8xcos(2x)-4x^2 sin(2x)-12x}{20x^3},\frac{0}{0}\] \[\frac{-6sin^2(x)-6xsin(2x)-2x^2cos(2x)}{15x^2},\frac{0}{0}\] \[\frac{-6sin(2x)-8xcos(2x)+2x^2sin(2x)}{15x},\frac{0}{0}\] \[\frac{-20cos(2x)+20xsin(2x)+4x^2cos(2x)}{15},x=0,-\frac{20}{15}=-\frac{4}{3}\]

OpenStudy (amistre64):

which aint 0, but of a = 4/3 then it might just work. I cant trust the wolf with this since it has no idea what to do with it :)

OpenStudy (amistre64):

it shows its good on the graph

OpenStudy (amistre64):

the limit from the left matches the limit from the right ....

OpenStudy (amistre64):

but then, thats only 1 value of a and b, i gots no idea what it would be like for "all" of them